Can I also have help with this one...
solve for x and y for x-y=4 and x+y=4
I am having such a hard time solving the problem... please help

Answer by navytutor(7)

First solve for any of the variables in any of the equations and then plug it in back in the other equation. x=4+y, then (4+y)+y=4 ---> 4+2y=4 2y=0 y=0, then x=4!!!